# Comprehensive Analysis of Relative Pressure Estimation Methods Utilizing 4D Flow MRI

### v1

@@ 4.3 Effect of Noise
In agreement with Bertoglio et al. [14], the average errors across all noisy realizations were roughly equivalent to the cases with no added noise, experimentally confirming that vWERP, STE, and PPE are unbiased estimators. Whereas higher resolution was almost always better in the SNR=inf cases, this trend did not precisely hold in the noisy cases. Low spatial resolution combined with high temporal resolution resulted in a higher spread in estimated pressures, although the average pressures across noise realizations were still close to the SNR=inf cases. Deviations from the baseline pressure drop estimation were also routinely higher in the TL across both in silico and in vitro analyses.
@@ 4.4 Pros and Cons of Pressure Estimators
Each pressure estimator recovered flow phantom ground truth pressures with a minimum correlation coefficient of 0.9 and a tendency to mildly underestimate peak pressure drops, albeit with varying severity. More clear differences in accuracy were present, however, in the sensitivity and spatiotemporal analyses. Based on our analysis, vWERP should be the preferred estimator when calculating one-dimensional pressure drops. It had the highest accuracy in the 4D flow MRI acquisitions and was significantly faster than STE. Although vWERP performed worse than PPE and STE in the in silico cases, vWERP’s improved performance in the in vitro cases suggests that there are aspects of MRI simulation that we are missing in our synthetic MRI generation methodology. When a fully 3D pressure field is desired, the ideal estimator depends on the SNR and resolution of the 4D flow acquisition. STE should be the standard choice (especially at lower resolutions). However, if a low SNR is expected, PPE may be better than STE. Similar to [15], STE outperformed PPE in terms of accuracy; however, PPE appeared to have a slight edge in its resistance to noise. PPE’s additional resistance to noise may be due to the slightly different implementation of PPE, which was implemented as an FVM rather than an FDM. It should be noted that Nolte et al. [15] used a different implementation of PPE and STE, utilizing backward differences for temporal derivatives (rather than midpoint) and an FEM. Although PPE was faster than STE (roughly 2.5 minutes vs. 8.7 minutes of CPU time), both methods calculated pressure fields in a reasonable amount of time.
@@ 5 Conclusion
In this work, we analyzed vWERP, STE, and PPE over in silico, in vitro, and in vivo datasets. Of the fully spatial estimators, STE was shown to be more accurate overall, whereas PPE was shown to be more resilient to noise. vWERP did not perform as well as either STE or PPE in the in silico analysis. In the in vitro analysis, STE continued to outperform PPE in accuracy but was mildly less resilient to corruption of the velocity field. vWERP moderately outperformed both STE and PPE in accuracy. PPE remained the most resilient to noise. Finally, the pressure estimators were evaluated qualitatively and for agreement over patient cases. PPE and STE agreed strongly, and vWERP tended to estimate the largest pressure drop, mimicking the trend of the in vitro analysis. All pressure estimators were remarkably resilient to noise and successfully estimated the ground truth pressure, albeit with tendencies to underestimate. The time to compute the pressure estimations was on the order of minutes, demonstrating the possibility of being integrated into a clinical workflow.

### v2

@@ 4.3 Effect of Noise
In agreement with Bertoglio et al. [14], the average errors across all noisy realizations were roughly equivalent to the cases with no added noise, experimentally confirming that vWERP, STE, and PPE are unbiased estimators. Whereas higher resolution was almost always better in the SNR=inf cases, this trend did not precisely hold in the noisy cases. Low spatial resolution combined with high temporal resolution resulted in larger deviations in estimated pressures, although the average pressures across noise realizations were still close to the SNR=inf cases. Deviation from the baseline relative pressure drop were also routinely higher in the TL compared to the FL across both in silico and in vitro analyses.
@@ 4.4 Pros and Cons of Pressure Estimators
Each pressure estimator recovered flow phantom ground truth pressures with a minimum correlation coefficient of 0.9 and a tendency to mildly underestimate peak pressure drops, albeit with varying severity. More clear differences in accuracy were present, however, in the sensitivity and spatiotemporal analyses. Based on our analysis, vWERP outperformed STE and PPE when calculating one-dimensional pressure drops. It had the highest accuracy in the 4D-flow MRI acquisitions and was significantly faster than STE. Although vWERP performed worse than PPE and STE in the in silico cases, vWERP’s improved performance in the in vitro cases suggests that there are aspects of MRI simulation that we are missing in our synthetic MRI generation methodology. When a fully 3D pressure field is desired, the ideal estimator depends on the SNR and resolution of the 4D-flow acquisition. Our data suggests that STE performs better, especially at lower spatial resolutions. However, if a low SNR is expected, PPE may outperform STE. Similar to [15], STE outperformed PPE in terms of accuracy; however, PPE appeared to have a slight edge in its resistance to noise. PPE’s additional resistance to noise may be due to the slightly different implementation of PPE, which was implemented as an FVM rather than an FDM. It should be noted that Nolte et al. [15] used a different implementation of PPE and STE, utilizing backward differences for temporal derivatives (rather than midpoint) and an FEM. Although PPE was faster than STE (roughly 2.5 minutes vs. 8.7 minutes of CPU time), both methods calculated pressure fields in a reasonable amount of time.
@@ 5 Conclusion
In this work, we analyzed vWERP, STE, and PPE over in silico, in vitro, and in vivo datasets. Of the fully spatial estimators, STE was shown to be more accurate overall, whereas PPE was shown to be more resilient to noise. vWERP did not perform as well as either STE or PPE in the in silico analysis. In the in vitro analysis, STE continued to outperform PPE in accuracy but was mildly less resilient to corruption of the velocity field. vWERP moderately outperformed both STE and PPE in accuracy. PPE remained the most resilient to noise. Finally, the pressure estimators were evaluated qualitatively and for agreement over patient cases. PPE and STE agreed strongly, and vWERP tended to estimate the largest pressure drop, mimicking the trend of the in vitro analysis. All pressure estimators were remarkably resilient to noise and successfully estimated the ground truth pressure, albeit with a tendency to underestimate peak relative pressures. The time to compute the pressure estimations was on the order of minutes, demonstrating the possibility of being integrated into a clinical workflow.